OpenStudy (anonymous):

9^(3x)=81^(x-4) All I got is 9^(3x)=(9^2)^(x-4) What do i do next?

jimthompson5910 (jim_thompson5910):

Multiply the exponents 9^(3x)=(9^2)^(x-4) 9^(3x) = 9^(2(x-4)) since the bases are equal, the exponents must be equal, so... 3x = 2(x-4)

jimthompson5910 (jim_thompson5910):

solve that for x to find your answer

OpenStudy (anonymous):

x=-8?

jimthompson5910 (jim_thompson5910):

you are correct

OpenStudy (anonymous):

THX!!!

jimthompson5910 (jim_thompson5910):

you're welcome